# Symplectic torus actions on complete intersections

### v1

@@ 3.1 Proving the JR equation for $4$-dimensional fixed components
$[\omega]|_{N}=c_{1}(M)|_{N}=c_{1}(N)+(\frac{1}{|H_{\max}|}+\frac{1}{|H_{\min}|})[\omega]|_{N},$
By, Corollary 2.10 $b_{2}^{+}(N)=1$, proving the theorem.
@@
Consider $\mathbb{CP}^{4}$ with the Hamiltonian $S^{1}$-action.
$z.[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]=[x_{0}:zx_{1}:zx_{2}:zx_{3}:z^{2}x_{4}].$
Then $M_{\min}$ and $M_{\max}$ are isolated and are mapped to $-5$ and $5$ respectively by the moment map of Theorem 2.8. The only other fixed component is $N=\mathbb{CP}^{2}$ with $H(N)=0$ and normal bundle $\mathcal{O}(1)\oplus\mathcal{O}(1)$. Therefore $c_{1}(M)|N=c_{1}(T\mathbb{CP}^{2})+2c_{1}(\mathcal{O}(1))=5H$, showing that $c_{1}(M)|N=|H_{\max}|c_{1}(\mathcal{O}(1))=|H_{\min}|c_{1}(\mathcal{O}(1))$.
We also present the following corollaries which may be of independent interest.
@@
Suppose that $M$ is an $8$-dimensional closed symplectic manifold with a Hamiltonian circle action and with $b_{2}(M)=1$. Then,
@@ 4 Results in arbitrary dimensions
4.0.1 Symplectic circle actions
In the following theorem, we give a classification of the multidegrees of complete intersections admitting a Hamiltonian $S^{1}$-action whose fixed point sets satisfy an analogous assumption to [20, Theorem 1.1].
@@ 4.0.1 Symplectic circle actions
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic circle action with no fixed components $F$ satisfying $\dim(F)>0$ and $\dim(M)=\dim(F)\mod 4$.
Then $M$ is diffeomorphic to $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ or $X_{2}(3),$ $m\in\mathbb{N}$.
@@
Firstly, suppose that $\dim(M)=2n$, with $n$ odd. By applying the Lefschetz hyperplane theorem and Poincaré duality to $M$, all of the even Betti numbers of $M$ are equal to $1$ and all of the odd Betti numbers are 00, except possibly $b_{n}(M)$.
Let $F$ be a fixed submanifold, then by Theorem 2.2, $b_{i}(F)\neq 0$ for at most one odd $i$. Since, by assumption $\dim(F)$ is divisible by $4$ this implies by applying Poincaré duality to $F$ that all of the odd Betti numbers of $F$ vanish, because every odd Betti number appears in at least two degrees. Therefore by Theorem 2.2 the odd Betti numbers of $M$ vanish. Since all of the even Betti numbers are equal to one, it follows that $\chi(M)=n+1$. Hence, by Theorem 2.11 the multidegree of $M$ is (1) or (2).
In the case that $\dim(M)$ is divisible by $4$, then by our assumptions all terms on the right hand side of the equation given in Theorem 3.2 vanish. It follows that
$\sigma(M)=\sum_{k\in\mathbb{Z}}(b_{4k}(M)-b_{4k+2}(M)).$
By Theorem 2.12 the multidegree of $M$ is $(1),(2)$ or $(2,2)$, or $M$ is a cubic surface. ∎
4.0.2 GKM actions
In this section, we strengthen the above results under the additional assumption that the action is GKM and preserves a symplectic form homotopic to the Fubini-Study form. We are able to give a complete characterization as in the classification of infinite automorphism groups due to Benoist [1, Theorem 3.1].
@@ 4.0.2 GKM actions
Let $X$ be a smooth complete intersection of complex dimension greater than $2$, and let $\omega_{0}$ be the restriction of the Fubini-Study form to $X$. Then $X$ admits a GKM torus action preserving a symplectic form that is homotopic to $\omega_{0}$ $\iff$ $|Aut(X)|=\infty$.
@@
By Theorem 1.2 and the theorem of Benoist [1, Theorem 3.1], it is sufficient to exclude the existence of such actions on a complete intersection of the form $X_{2m}(2,2)$. Suppose that there is such a GKM action. Set $n=2m$ to be the complex dimension of $X_{2m}(2,2)$. Since $X_{2m}(2,2))$ is Fano it is possible to rescale the hypothetical invariant symplectic form so that $c_{1}=[\omega]$.
A calculation with the adjunction formula and Bezout’s theorem gives that the the Chern number $c_{1}c_{n-1}$ of $X_{n}(2,2)$ is given by
$\frac{n}{2}(n+2)(n-1).$
Moreover, the topological Euler characteristic is
@@
The collection of spheres corresponding to the edges is Poincaré dual to $c_{n-1}$ [14] and by the adjunction formula $c_{1}$ is $(n-1)$ times the primitive generator of $H^{2}$. Therefore, using the condition $c_{1}=[\omega]$ the integral of $c_{1}$ on each of the spheres is a positive integer multiple of $(n-1)$. Therefore implying that the Chern number $c_{1}c_{n-1}$ is divisible by $n(n+2)(n-1)$, which is a contradiction.
We note finally, that it is possible to obtain a statement covering complex dimension $2$, at the price of requiring that the symplectic form preserved by the GKM action is monotone. Since both $X_{2}(2,2)$ and $X_{2}(3)$ admit Kähler forms preserved by a toric action, this is the strongest possible statement which can cover complex dimension $2$. Since the proof is identical to the proof of Theorem 4.2, it is omitted.
@@
Another, perhaps more elegant but weaker possibility, is to require that the symplectic structure is symplectomorphic to one given by the Fubini-Study symplectic form. This was stated in Theorem 1.4 of the introduction.
@@ 5 Unimodality of Betti numbers
Given the generality of Proposition 3.2, it is natural ask whether $I_{JR}(M)=0$ for all closed symplectic manifolds admitting a Hamiltonian circle action. The following example shows that this is not the case.
Consider the standard Hamiltonian $S^{1}$-action on $\mathbb{CP}^{4}$ fixing a point and a hyperplane $H$. Let $S\subset H$ be a smooth degree $4$ hypersurface. Recall that $S$ is a K3 surface with signature $16$ and $b_{2}(S)=22$. Let $M$ be the $S^{1}$-equivariant blow up of $\mathbb{CP}^{4}$ along $S$. Then by Proposition 3.2
$I_{JR}(M)=-I_{JR}(N)=-36.$
In this section, we will present some applications regarding the unimodality of Betti numbers of manifolds admitting a Hamiltonian circle action. Tolman has asked whether for a closed symplectic manifold with a Hamiltonian circle action, the even Betti numbers are unimodal (i.e. weakly increasing up to the middle degree) [15, Problem 4.3]. For Kähler manifolds the property holds from the hard Lefschetz theorem.
In [5] Cho and Kim proved that $8$-dimensional closed symplectic manifolds having a Hamiltonian $S^{1}$-action with isolated fixed points, the even Betti numbers are unimodal. Here we prove the same result under the weaker assumption that neither of the extremal submanifolds are $4$-dimensional.
Theorem 5.1.
Suppose that $(M,\omega)$ is a closed symplectic $8$-manifold admitting a Hamiltonian circle action such that, $M_{\min},M_{\max}$ are not $4$-dimensional. Then the even Betti numbers of $M$ are unimodal.
Proof.
By Theorem 3.2 $M$ satisfies:
@@
Since by assumption $M_{\min}$ and $M_{\max}$ are not $4$-dimensional, every $4$-dimensional fixed component $F$, satisfies $\lambda_{F}=1$.
Therefore the above equation becomes:
@@
By Lemma 3.3 $I_{JR}(F)=2b^{+}(F)-2$, since $F$ is a symplectic $4$-manifold, this gives $I_{JR}(F)>0$. Therefore giving $I_{JR}(M)<0$.
Or, the inequality may me restated as
$\sigma(M)-2+2b_{2}(M)-b_{4}(M)<0.$
Then, since $M$ is symplectic $-b_{4}(M)<\sigma(M)$, substituting this and rearranging gives $1+b_{2}(M)<b_{4}(M)$. Then, since both $b_{2}(M),b_{4}(M)$ are integers this gives the desired inequality.
@@
Suppose that $M$ is a symplectically rational $8$-manifold admitting a Hamiltonian $S^{1}$-action. Then, the even Betti numbers of $M$ are unimodal.
If a complex rational $4$-fold $M$ has such an algebraic torus action having an algebraic surface $N$ as an extremal component, then $M$ is birational to a weighted projective space bundle over $N$ via a standard GIT argument. Therefore $N$ is stably rational and hence rational [30]. Therefore $N$ contributes 00 to the equation of Proposition 3.2 and the proof of Theorem 5.1 goes through.
We also obtain a similar result for closed $12$-manifolds with $b_{2}=1$. To the authors knowledge this result is new, under the weaker assumption of isolated fixed points.
@@
Suppose that $(M,\omega)$ is a closed symplectic $12$-manifold admitting a Hamiltonian circle action with no $4$ or $8$ dimensional fixed components, and $b_{2}(M)=1$. Then the even Betti numbers of $M$ are unimodal.
@@
By Theorem 3.2 $M$ satisfies the JR equation. Since the modulus of the signature of a symmetric bilinear form is bounded above by the rank of the middle homology group, we have that
$\sigma(M)\leq b_{6}(M).$
Then substituting the JR equation for $\sigma(M)$ gives
$2b_{4}(M)-b_{6}(M)\leq b_{6}(M).$
Which implies that $b_{4}(M)\leq b_{6}(M)$, as required. ∎
Next, we prove two inequalities on the Betti numbers of a closed symplectic manifold admitting a Hamiltonian $S^{1}$-action, under the assumption that there is no fixed components of positive dimension divisible by $4$. Under the assumption of isolated points, the inequalities were proved by Cho [3, Theorem 1.2]. Both inequalities may be considered as evidence for the unimodality of even Betti numbers of manifolds admitting a Hamiltonian $S^{1}$-action, in the sense that they would be implied by the unimodality. In particular, they imply that the possibilities for any particular Betti number of the form $b_{k}$ with $k=2n\mod 4$ are bounded, if all of the other Betti numbers are fixed, provided $k\neq n$.
@@
By Proposition 3.2 the signature of $M$ is equal to the alternating sum of its even degree Betti numbers.
The first inequality, follows $-b_{n}(M)<\sigma(M)$, after putting all terms to the left hand side, applying Poincaré duality and dividing by $2$. In this case the inequality is strict because the equality case is equivalent to the symplectic form being negative definite, and the top power of the symplectic form is positive.
The second inequality then follows from the inequality $\sigma(M)\leq b_{n}(M)$, where $n=\frac{\dim(M)}{2}$, after putting all terms to the right hand side, applying Poincaré duality and dividing by $2$.
We finally state a proposition regarding the Betti numbers of spin manifolds admitting a Hamiltonian circle action.
@@
Let $(M,\omega)$ be a closed symplectic manifold of dimension $2n$ with a Hamiltonian $S^{1}$-action having no fixed component of positive dimension divisible by $4$. Suppose that $n=2\mod 4$ and $M$ is spin, and $b_{2i}(M)=1$ for $2i\neq n$. Then $b_{n}(M)=2\mod 16$.

### v2

@@ 3.1 Proving the JR equation for $4$-dimensional fixed components
$\omega]|_{N}=c_{1}(M)|_{N}=c_{1}(N)+(\frac{1}{|H_{\max}|}+\frac{1}{|H_{\min}|})[\omega]|_{N},$
Using the fact that $|H_{\min}|\geq 4$ and $|H_{\min}|\geq 4$ implies that
$1-(\frac{1}{|H_{\max}|}+\frac{1}{|H_{\min}|})>0.$
By, Corollary 2.12 $b_{2}^{+}(N)=1$, proving the theorem.∎
@@
Consider $\mathbb{CP}^{4}$, with the Fubini-Study symplectic form, rescaled so that $c_{1}=[\omega]$. Consider the Hamiltonian $S^{1}$-action, defined for $\lambda\in S^{1}$ by:
$\lambda\cdot[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]=[x_{0}:\lambda x_{1}:\lambda x_{2}:\lambda x_{3}:\lambda^{2}x_{4}].$
Let $X_{h}\in H^{2}(\mathbb{CP}^{4},\mathbb{Z})$ denote the hyperplane class i.e. the Poincaré dual of a hyperplane  [^3] We deviate from the usual notation $H$ because $H$ denotes the Hamiltonian.. For the above action, $M_{\min}$ and $M_{\max}$ are isolated and are mapped to 5-5 and $5$ respectively by the moment map of Theorem 2.10. The only other fixed component is $N=\mathbb{CP}^{2}$ with $H(N)=0$ and normal bundle $\mathcal{O}(1)\oplus\mathcal{O}(1)$. Therefore $c_{1}(M)|N=c_{1}(T\mathbb{CP}^{2})+2c_{1}(\mathcal{O}(1))=5X_{h}$, showing that $c_{1}(M)|N=|H_{\max}|c_{1}(\mathcal{O}(1))=|H_{\min}|c_{1}(\mathcal{O}(1))$.
Next, we show that Theorem 1.2 follows from Theorem 3.6.
Proof of Theorem 1.2.
Suppose that $(M,\omega)$ is an $8$-dimensional complete intersection with a symplectic $S^{1}$-action. Because $M$ is simply connected the action is Hamiltonian. By Lemma 2.14 $b_{2}(M)=1$. By Theorem 3.6 the intersection form of $M$ is positive definite. Therefore by Theorem 2.16 $M$ is diffeomorphic to either $X_{4}(1),X_{4}(2)$ or $X_{4}(2,2)$.∎
We also present the following corollary which may be of independent interest.
@@
Suppose that $M$ is an $8$-dimensional closed symplectic manifold with a Hamiltonian $S^{1}$-action and with $b_{2}(M)=1$. Then,
@@ 4 Results in arbitrary dimensions
4.1 Symplectic $S^{1}$-actions
In the following theorem, we give a classification of the multidegrees of complete intersections admitting a Hamiltonian $S^{1}$-action whose fixed point sets satisfy an assumption analogous to that of [29, Theorem 1.1].
@@ 4.1 Symplectic $S^{1}$-actions
Suppose that $(M,\omega)$ is a complete intersection admitting a symplectic $S^{1}$-action with no fixed components $F$ satisfying $\dim(F)>0$ and $\dim(M)=\dim(F)\mod 4$. Then $M$ is diffeomorphic to one of $X_{m}(1),X_{m}(2),X_{2m}(2,2),$ for some $m\in\mathbb{N}$ or $X_{1}(3)$, $X_{2}(3)$.
@@
If $\dim(M)=2$ by the classification of smooth $S^{1}$-actions on closed orientable surfaces the statement holds. Therefore assume $\dim(M)=2n>2$, then since $\pi_{1}(M)=\{1\}$ by the Lefschetz hyperplane theorem, the action is Hamiltonian.
Firstly, suppose that $\dim(M)=2n$, with $n$ odd. By Lemma 2.14 all of the even Betti numbers of $M$ are equal to $1$ and all of the odd Betti numbers are 0, except possibly $b_{n}(M)$. Let $F$ be a fixed submanifold. By Theorem 2.2, $b_{i}(F)\neq 0$ for at most one odd $i$. Since by assumption $\dim(F)$ is divisible by $4$, this implies by applying Poincaré duality to $F$ that all of the odd Betti numbers of $F$ vanish, because every odd Betti number appears in at least two degrees. Therefore by Theorem 2.2 the odd Betti numbers of $M$ vanish. Since all of the even Betti numbers are equal to one, it follows that $\chi(M)=n+1$, hence the result follows by Theorem 2.15.
Next, suppose that $\dim(M)=2n$, with $n$ even. Then by assumption all terms on the right hand side of the equation given in Equation (1) of Proposition 3.2 vanish. It follows that
$\sigma(M)=\sum_{k\in\mathbb{Z}}(b_{4k}(M)-b_{4k+2}(M)),$
hence the result follows by Theorem 2.16.∎
4.2 GKM actions
In this section, we strengthen Theorem 4.1 under the additional assumption that the action is GKM and preserves a symplectic form homotopic to the Fubini-Study form. We are able to give a complete characterization as in the classification of infinite automorphism groups due to Benoist (see Theorem 2.18).
@@ 4.2 GKM actions
Let $X$ be a smooth complete intersection of complex dimension greater than $2$, and let $\omega_{FS}$ be the restriction of the Fubini-Study form to $X$. Then $X$ admits a GKM torus action preserving a symplectic form $\omega$ that is homotopic to $\omega_{FS}$ $\iff$ $|Aut(X)|=\infty$.
@@
By Corollary 1.3 and the theorem of Benoist (see Theorem 2.18), it is sufficient to exclude the existence of such an action on $X_{2m}(2,2)$ for some positive integer $m$, set $n=2m$.
Suppose that there is a closed symplectic manifold $(M,\omega)$ diffeomorphic to $X_{n}(2,2)$ such that $\omega$ is homotopic to $\omega_{FS}$ and such that $(M,\omega)$ has a Hamiltonian, GKM torus action. By Lemma 2.14 and [28, Proposition 3.4]:
[…]
Applying [14, Corollary 3.1] gives
$\int_{M}c_{1}c_{n-1}(M)=\sum_{p=0}^{n}b_{2p}(M)(6p(p-1)+\frac{5n-3n^{2}}{2}).$
Inputting Equation (3) and simplifying gives
$\int_{M}c_{1}c_{n-1}(M)=\frac{n}{2}(n+2)(n-1).$
Since by assumption $n>2$ it follows by Lemma 2.14 that $b_{2}(M)=1$. Hence by Lemma 2.9 it is possible to rescale the hypothetical invariant symplectic form so that $c_{1}=[\omega]$.Moreover, the topological Euler characteristic of $M$ is
@@
By Lemma 2.5, the collection of spheres corresponding to the edges is Poincaré dual to $c_{n-1}$. By the adjunction formula $c_{1}$ is (n1)(n-1) times the primitive generator of $H^{2}(X_{n}(2,2))\cong\mathbb{Z}$. Therefore, using the condition $c_{1}=[\omega]$ the integral of $c_{1}$ on each of the spheres is a positive integer multiple of (n1)(n-1). Hence the Chern number $c_{1}c_{n-1}$ is bounded below by n(n+2)(n1)n(n+2)(n-1) which is a contradiction. ∎
We note finally that it is possible to obtain a statement covering complex dimension $2$, at the price of requiring that the symplectic form preserved by the GKM action is monotone. Since both $X_{2}(2,2)$ and $X_{2}(3)$ admit Kähler forms preserved by a toric action, this is the strongest possible statement which can cover complex dimension $2$. Since the proof is identical to the proof of Theorem 4.2, it is omitted.
@@
Remark 4.4.
Another, perhaps more elegant but weaker, possibility is to require that the symplectic structure is symplectomorphic to one given by the Fubini-Study symplectic form. This was stated in Theorem 1.5 of the introduction.
@@ 5 Unimodality of Betti numbers
In this section, we will present some applications regarding the unimodality of Betti numbers of manifolds admitting a Hamiltonian $S^{1}$-action. Tolman has asked whether for a closed symplectic manifold with a Hamiltonian $S^{1}$-action, the even Betti numbers are unimodal (i.e. weakly increasing up to the middle degree) [22, Problem 4.3]. For Kähler manifolds the property follows from the hard Lefschetz theorem.
In [6] Cho and Kim proved that $8$-dimensional closed symplectic manifolds having a Hamiltonian $S^{1}$-action with isolated fixed points, the even Betti numbers are unimodal. Here we prove the same result under the weaker assumption that neither of the extremal submanifolds is $4$-dimensional.
Proof of Theorem 1.6.
By Proposition 3.2,
@@
Since by assumption $M_{\min}$ and $M_{\max}$ are not $4$-dimensional, every $4$-dimensional fixed component $F$ satisfies $\lambda_{F}=1$.
@@
By Lemma 3.5 and since $F$ is a closed symplectic submanifold:
$I_{JR}(F)=2b^{+}(F)-2\geq 0,$
hence by Poincaré duality
$0\geq I_{JR}(M)=\sigma(M)-2+2b_{2}(M)-b_{4}(M).$
Since $M$ is a closed symplectic manifold $\sigma(M)>-b_{4}(M)$. Hence
$b_{2}(M)<b_{4}(M)+1,$
and so the result follows. ∎
@@
If $(M,\omega)$ is a closed symplectically rational $8$-manifold admitting a Hamiltonian $S^{1}$-action. Then, the even Betti numbers of $M$ are unimodal.
If a complex, projective and rational $4$-fold $M$ has such an algebraic torus action having an algebraic surface $N$ as an extremal component, then $M$ is birational to a weighted projective space bundle over $N$ via a standard GIT argument. Therefore $N$ is stably rational and hence rational [40]. Hence $N$ contributes 0 to the equation of Proposition 3.2 and the proof of Theorem 1.6 goes through.
We also obtain a similar result to Theorem 1.6 for closed $12$-manifolds with second Betti number equal to $1$. To the authors knowledge this result is new, under the stronger assumption of isolated fixed points.
@@
Suppose that $(M,\omega)$ is a closed symplectic $12$-manifold such that $b_{2}(M)=1$. Suppose that $M$ admits a Hamiltonian $S^{1}$-action with no 44- or 88-dimensional fixed components, then the even Betti numbers of $M$ are unimodal.
@@
By Proposition 3.2 we have that $I_{JR}(M)=0$. Since $\sigma(M)\leq b_{6}(M),$ by Poincaré duality,
$I_{JR}(M)=0\;\implies\;2b_{4}(M)-b_{6}(M)\leq b_{6}(M),$
and so the result follows.∎
Next we prove two inequalities on the Betti numbers of a closed symplectic manifold admitting a Hamiltonian $S^{1}$-action, under the assumption that there is no fixed components of positive dimension divisible by $4$. Under the assumption of isolated fixed points, the inequalities were proved by Cho [4, Theorem 1.2]. Both inequalities may be considered as evidence for the unimodality of even Betti numbers of manifolds admitting a Hamiltonian $S^{1}$-action, in the sense that they would be implied by the unimodality. We formulate this as a remark after Proposition 5.3
@@
By assumption all terms on the right hand side of the equation given in Equation (1) of Proposition 3.2 vanish. Thus by Proposition 3.2 $M$ satisfies the JR equation.
The first inequality follows from the the inequality $b_{n}(M)<\sigma(M)$, after putting all terms to the left hand side, applying Poincaré duality and dividing by $2$. In this case the inequality is strict because the equality case is equivalent to the symplectic form being negative definite, and the top power of the symplectic form is positive.
The second inequality then follows from the inequality $\sigma(M)\leq b_{n}(M)$, after applying Poincaré duality and dividing by $2$. ∎
Remark 5.4.
Suppose that $(M,\omega)$ is a closed symplectic $2n$-manifold and the even Betti numbers are unimodal. Then the inequalities of Proposition 5.3 hold.
To see this, suppose that $4|n$ and $b_{2i}(M)\leq b_{2i+2}(M)$ for all integers $i$ such that $2i+2\leq n$. Then rewriting the quantity $1-b_{2}+b_{4}-\ldots+b_{n}$, as
$1+(-b_{2}+b_{4})+\ldots+(-b_{n-2}+b_{n}),$
shows that it is manifestly positive. The case $n=2\mod 4$ is proved similarly (as in Proposition 5.3 the inequality is not strict in this case). We remark also that the inequalities of Proposition 5.3 imply that the possibilities for any particular Betti number of the form $b_{k}$ with $k=2n\mod 4$ are bounded, if all of the other Betti numbers are fixed, provided $k\neq n$.
We finally state a proposition regarding the Betti numbers of spin manifolds admitting a Hamiltonian $S^{1}$-action.
@@
Let $(M,\omega)$ be a closed symplectic manifold of dimension $2n$ with a Hamiltonian $S^{1}$-action having no fixed component of positive dimension divisible by $4$. Suppose that $n=2\mod 4$, that $M$ is spin, and that $b_{2i}(M)=1$ for $2i\neq n$. Then $b_{n}(M)=2\mod 16$.